OpenStudy (anonymous):

The length of a rectangle is 7ft less than 2 times its width. find the dimensions of the rectangle if the area is 72 sq ft

OpenStudy (anonymous):

please help word problems just confuse me

OpenStudy (anonymous):

let the width of of rectangle be x so if width is x then acrding ot ur question "The length of a rectangle is 7ft less than 2 times its width" it => lenght = 2x -7 now u can do it

OpenStudy (anonymous):

A=L*W L= 2W-7 72 = (W)*(2W-7) can u solve from now?

OpenStudy (anonymous):

time and move all in one side , factor

OpenStudy (anonymous):

thanks so much!